OpenStudy (anonymous):

(4+i)^2 - (6-i)^2 write result in standard form a+bi

jimthompson5910 (jim_thompson5910):

(4+i)^2 - (6-i)^2 16+8i+i^2 - (36-12i+i^2) 16+8i-1 - (36-12i-1) 16+8i-1 - 36+12i+1 -20+20i So (4+i)^2 - (6-i)^2 = -20+20i

jimthompson5910 (jim_thompson5910):

Keep in mind that \[\large i=\sqrt{-1}\] so \[\large i^2=-1\]